Consider the following information to validate the hypothesis that the mean of the population


is at most 30 at 5% level of significance.


Sample size = 45, mean of the sample = 38, variance = 4



What will be conclusion if the level of significance is taken as 1%? Mention the observations, if


any.



1
Expert's answer
2021-07-01T09:48:47-0400

1. The following null and alternative hypotheses need to be tested:

"H_0: \mu\leq 30"

"H_1:\mu>30"

This corresponds to a right-tailed test, for which a t-test for one mean, with unknown population standard deviation, using the sample standard deviation, will be used.


Based on the information provided, the significance level is "\alpha=0.05,"

"df=n-1=45-1=44" degrees of freedom, and the critical value for a right-tailed test is "t_c=1.68023."

The rejection region for this right-tailed test is "R=\{t:t>1.68023\}."

The t-statistic is computed as follows:


"t=\dfrac{\bar{x}-\mu}{s/\sqrt{n}}=\dfrac{38-30}{4/\sqrt{45}}=13.4164"

Since it is observed that "t=13.4164>1.68023=t_c," it is then concluded that the null hypothesis is rejected.

Using the P-value approach: The p-value for "\alpha=0.05, df=44, t=13.4164,"right-tailed is "p=0," and since "p=0<0.05=\alpha," it is concluded that the null hypothesis is rejected.


Therefore, there is not enough evidence to claim that the population mean "\mu"

"\mu"  is at most 30, at the "\alpha=0.05" significance level.



2. The following null and alternative hypotheses need to be tested:

"H_0: \mu\leq 30"

"H_1:\mu>30"

This corresponds to a right-tailed test, for which a t-test for one mean, with unknown population standard deviation, using the sample standard deviation, will be used.


Based on the information provided, the significance level is "\alpha=0.01,"

"df=n-1=45-1=44" degrees of freedom, and the critical value for a right-tailed test is "t_c=2.414134."

The rejection region for this right-tailed test is "R=\{t:t>2.414134\}."

The t-statistic is computed as follows:


"t=\dfrac{\bar{x}-\mu}{s/\sqrt{n}}=\dfrac{38-30}{4/\sqrt{45}}=13.4164"

Since it is observed that "t=13.4164>2.414134=t_c," it is then concluded that the null hypothesis is rejected.

Using the P-value approach: The p-value for "\alpha=0.05, df=44, t=13.4164,"right-tailed is "p=0," and since "p=0<0.01=\alpha," it is concluded that the null hypothesis is rejected.


Therefore, there is not enough evidence to claim that the population mean "\mu"

"\mu"  is at most 30, at the "\alpha=0.01" significance level.


Therefore there is not enough evidence to claim that the population mean "\mu"

"\mu"  is at most 30, at "\alpha=0.05" significance level and at "\alpha=0.01" significance level.